OpenStudy (anonymous):

Solve simultaneously for "x" and "y": 7x + 2y = -4 3x + 4y = 14

OpenStudy (anonymous):

http://www.twiddla.com/563490

OpenStudy (anonymous):

You have to come to twiddla and post your problem in order for me to help...

OpenStudy (anonymous):

\[7x+2y=-4\]\[3x+4y=14\]\[7x+2y=-4..................*3\]\[3x+4y=14................*7\]\[21x+6y=-12\]\[21x+28y=98\]\[6y-28y=-22y\]\[-12-98=-110\]\[-22y=-110\]\[y=-110/-22.........=5\] then substitute to get x

OpenStudy (anonymous):

Solving the matrix\[\left[\begin{matrix}7 & 2 & -4\\ 3 & 4 &14\end{matrix}\right]\] it's \[\left[\begin{matrix}1 & 0 & (-14/7)\\ 0 & 1 & 5\end{matrix}\right]\] and the answer is x= -14/7 and y= 5